OpenStudy (anonymous):

Can someone help me with this one? I would greatly appreciate it! 8x-3=4x+5

OpenStudy (anonymous):

are you solving for x?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok so you need to put like terms together...do you know what that means?

OpenStudy (anonymous):

Not really, no.

OpenStudy (ivancsc1996):

Putting all the x's in one side and the terms without on the other.

OpenStudy (anonymous):

ok so you need to put the xs together and the normal numbers together so subtract 4x from both sides

OpenStudy (anonymous):

because whatever you do on one side of the equals sign you have to do to the other. think of the equals sign as a mirror. every move you make one one side has to be reflected on the other side

OpenStudy (anonymous):

I see. So it would be 8x-4x-3=4x-4x+5?

OpenStudy (anonymous):

yes exactly and you do the same thing to the normal numbers

OpenStudy (anonymous):

So then after that it would be 4x-3=x+5? Or is it just +5 at the end?

OpenStudy (anonymous):

Or did I get it absolutely wrong?

OpenStudy (anonymous):

you are close but you need to move that 3 over with the 5 because they are like terms so to move the -3 over you need to ADD 3 to both sides

OpenStudy (anonymous):

but we're not done yet! to solve for x we need to get x all by itself

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

so what do you have right now?

OpenStudy (anonymous):

I'm not entirely sure, but on paper I have; 4x+3=x+5-3+3 I don't know if that's correct...

OpenStudy (anonymous):

ok so this is what we have done so far: 8x-3=4x+5 -4x -4x subtract 4x from both sides 4x-3=5 +3 +3 add three on both sides to get rid of the -3 so you should have 4x=8

OpenStudy (anonymous):

do you understand this so far?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok so to solve for x we need to get x all by itself right? So think of 4x as 4 times x now 4 times x is multiplication so to get x by itself we need to do the OPPOSITE of multiplication which is...?

OpenStudy (anonymous):

Division.

OpenStudy (anonymous):

yes! so divide 4 from both sides 4x=8 4/4 8/4

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

You'll have to excuse me if I'm being a bit slow or if I don't follow. English isn't my native language and I've never tried to have a conversation including mathematic terms... I get 0 and 2, so that means x=2, right?

OpenStudy (anonymous):

yes! very good but you should get 1 when you divide 4 by 4 so 1 times x =x but you wrote it right

OpenStudy (anonymous):

Ah, my bad.

OpenStudy (anonymous):

what is your native language?

OpenStudy (anonymous):

Norwegian.

OpenStudy (anonymous):

is that any close to german?

OpenStudy (anonymous):

They are related, yes.

OpenStudy (anonymous):

well your english is very good and if you need any more help just let me know

OpenStudy (anonymous):

Thank you, and thank you so much for your time, you really helped me alot!